OpenStudy (anonymous):

fill in the blank to make the expression a perfect square v^2+2v+___

OpenStudy (anonymous):

Do it with leters first: (a+b)^2=(a+b)(a+b)=a^2+ab+ab+b^2=a^2+2ab+b^2 now you already have a, just put 2v=2ab=2vb--->b=1 and (v+1)^2

OpenStudy (anonymous):

that didnt help...